Before lunch we took a closer look at two of the suspensions that were most problematic: ITMY and ETMY. Over lunch we took new free swinging data. Results below:
- For ITMY we discovered that the whitening on the UL sensor was not switching. This was causing the UL sensor to have a different response, with a steeper roll of, which was causing all of the transfer function estimates to the other sensors to have large imaginary components. We took new free swing data with all of the whitening turned OFF. The result is a much improved matrix and diagnalization. The input matrix elements are mostly the same, but the coupling is basically gone. We'll fix the whitening after the pump down.
ITMY |
 |
pit yaw pos side butt
UL 0.157 1.311 1.213 -0.090 0.956
UR 1.749 -0.490 0.886 -0.038 -1.042
LR -0.251 -2.000 0.787 -0.007 1.066
LL -1.843 -0.199 1.114 -0.059 -0.936
SD -0.973 -0.205 1.428 1.000 0.239 |
4.34779 |
- ETMY has a very problematic SIDE OSEM. The magnet does not line up with the OSEM axis, and since there is no lateral adjustment in the side OSEMs, there's not much we can do about this. We're using aluminum foil to wedge the OSEM over as far as possible, but it's not quite enough. With the OSEM plates horizontal there is a lot of POS->SIDE coupling. With the OSEM plates vertical, the magnetic sits a little too close to the rear face, which can cause the magnet to get stuck to the LED plate. We're trying to decide where to leave it now, but the new diagnalization with the OSEM plates vertical is definitely better:
ETMX |
 |
pit yaw pos side butt
UL -0.138 1.224 1.463 -0.086 0.944
UR 0.867 -0.776 1.501 -0.072 -1.051
LR -0.995 -0.896 0.537 -0.045 0.754
LL -2.000 1.104 0.499 -0.059 -1.251
SD 0.011 0.220 1.917 1.000 0.224 |
4.42482 |
|